# CFG++: Manifold-constrained Classifier Free Guidance for Diffusion Models

### v1

@@ 3.1 Derivation of CFG++
Instead of uncritically adopting the sharpened posterior distribution $p^{\omega}({\boldsymbol{x}}|{\boldsymbol{c}})\propto p({\boldsymbol{x}})p({\boldsymbol{c}}|{\boldsymbol{x}})^{\omega}$ as introduced by ho2021classifierfree and further explored in the Bayesian CFG formulation, we adopt a fundamentally different strategy by reformulating text-guidance as an optimization problem. Specifically, our focus is on identifying a loss function $\ell({\boldsymbol{x}})$ in (8) such that, when minimized under the condition set by the text, enables the reverse diffusion process to progressively generate samples that increasingly satisfy the text condition.
One of the most significant contributions of this paper is to reveal that the text-conditioned score matching loss or score distillation sampling (SDS) loss is also ideally suited for our purpose. Specifically, we are interested in solving the following inverse problem through diffusion models:
@@
This implies that our goal is to identify solutions on the clean manifold ${\mathcal{M}}$ that optimally align with the text condition ${\boldsymbol{c}}$.
To avoid the Jacobian computation, in this paper we attempt to solve (12) through decomposed diffusion sampling (DDS). The resulting sampling process from reverse diffusion is then given by
${\boldsymbol{x}}_{t-1}$ $=$ $\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell_{sds}({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}$ (13)
which is analogous to (11). Consequently, the remaining technical issue is to compute $\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell_{sds}({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})$. To answer this, by using the forward diffusion to generate ${\boldsymbol{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\boldsymbol{x}}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}$ from the clean image ${\boldsymbol{x}}\in{\mathcal{M}}$, we convert the objective function in (12) into the following equivalent loss:
$\ell_{sds}({\boldsymbol{x}})$ $=\|{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})\|^{2}=\left\|\frac{{\boldsymbol{x}}_{t}-\sqrt{\bar{\alpha}_{t}}{\boldsymbol{x}}}{\sqrt{1-\bar{\alpha}_{t}}}-\frac{{\boldsymbol{x}}_{t}-\sqrt{\bar{\alpha}_{t}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}}{\sqrt{1-\bar{\alpha}_{t}}}\right\|^{2}$ (14)
$=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}\|{\boldsymbol{x}}-{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}\|^{2}$ (15)
where we use the conditional Tweedie formula derived by replacing ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}$ in (4) by the conditioned score ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}$. Similar observation has been made in DreamSampler (kim2024dreamsampler) and HiFA (zhu2023hifa). Consequently, $\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell_{sds}({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})$ can be readily computable and the resulting DDS sampling process can be simplified as
@@
where $\lambda:=\frac{2{\bar{\alpha}_{t}}}{{1-\bar{\alpha}_{t}}}\gamma_{t}$. Although concise, (16) can be further simplified to a similar form of the traditional CFG as described in Algorithm 1. Specifically, note that
${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}+\lambda({\hat{\boldsymbol{x}}_{\boldsymbol{c}}}-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})=({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$ (17)
where
${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}):={\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+{\lambda}[{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})]$ (18)
Therefore, (16) can be equivalently represented by:
@@
which is summarized in Algorithm 2.
By examining Algorithm 1 and Algorithm 2, we observe a notable similarity between CFG and CFG++. However, a crucial difference exists. Specifically, in CFG++, the renoising process after applying Tweedie’s formula should utilize the unconditional noise ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ instead of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$. This surprisingly simple fix to the original CFG algorithm leads to smoother trajectory of generation, (Fig. 1 bottom) and generation with superior quality (Fig. 1 top). Moreover, CFG++ can also be easily extended to significantly improve DDIM inversion, which is known to be prone to failure by standard CFG (mokady2023null). See Fig. 1 (middle) for representative results and Sec. 4.2 for further discussions. In the following section, we present a theoretical analysis of CFG++ to investigate the origin of this improvement.
@@ 3.2 Geometry of CFG++
Mitigating off-manifold phenomenon. In Fig. 1 (bottom), we illustrate the evolution of the posterior mean through Tweedie’s formula during the reverse diffusion process. Notably, in the early phases of reverse diffusion sampling under CFG, there is a sudden shift in the image and intense color saturation, followed by only minor changes until later stages. This abrupt transition in the posterior mean is somewhat unexpected, as the reverse diffusion process is designed to progressively refine resolution, moving from a noisier to a clearer manifold. Similar phenomena are depicted in Fig. 13 in Appendix. We refer to this as the off-manifold phenomenon, which will be clarified shortly. Conversely, under CFG++, the posterior mean evolves gradually, transitioning smoothly from a low-resolution image to high-resolution details. This indicates that CFG++ is less prone to the off-manifold phenomenon observed with CFG. In the following, we provide a theoretical analysis of the origin of the off-manifold phenomenon and explain why CFG++ can mitigate this issue.
By examining (17) and (19) as well as Algorithm 1, we can easily see that the Tweedie’s formula for the CFG++ and CFG can be equivalently represented by
@@
${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})=(1-\lambda){\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\lambda{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t}),\quad{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})=(1-\omega){\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\omega{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})$ (21)
@@
This implies that $\lambda,\omega\in[0,1]$ faciliates an interpolation between the unconditional and the conditional posterior means. However, given the typical CFG scale $\omega>1.0$, we observe that CFG extrapolates beyond the unconditional and the conditional posterior means established by Tweedie’s formula. Consequently, under the usual assumption that the clean manifold to be piecewise linear around the posterior mean derived via Tweedie’s formula, the conditional posterior mean estimates from traditional CFG, obtained with a guidance scale outside the range of $[0,1]$, can extend beyond the piecewise linear manifold. This may lead to the estimates potentially ‘falling off’ the data manifold, as depicted by an orange arrow pointing downwards in Fig. 3(a). We conjecture that this is one of the primary contributors to the off-manifold phenomenon. Thus, we select $\lambda\in[0,1]$ as the guidance scale for CFG++ to ensure it remains an interpolation between the unconditional and conditional Tweedie, thus preventing it from ‘falling off’ the clean data manifold.
An additional source of the off-manifold phenomenon in CFG occurs during the transition from the clean manifold ${\mathcal{M}}$ to the subsequent noisy manifold ${\mathcal{M}}_{t-1}$. Specifically, even from the same point in the clean manifold, the use of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ for this transition introduces a nonzero offset $\Delta^{\omega}:=\omega\sqrt{1-\bar{\alpha}_{t-1}}({\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t}))$ from ${\mathcal{M}}_{t-1}$, as illustrated in Fig. 3 (b). This leads to the invalid lifts from ${\mathcal{M}}$ to ${\mathcal{M}}_{t-1}$. On the other hand, we note that the renoising process of CFG++ follows the standard DDIM, resulting in the accurate transition to ${\mathcal{M}}_{t-1}$.
@@
Figure 5: The normalized text-conditioned score matching loss throughout the reverse diffusion sampling process for both CFG and CFG++.
Text-Image alignment. In Fig. 1 (top), we display images generated by SDXL using the prompt “Kayak in the water, optical color, areal view, rainbow". The generated image on the left is guided by the traditional CFG, while the one on the right is produced with our CFG++. Although both images look impressive and appear to align well with the given text prompt, a closer examination reveals a significant difference. For instance, one of the oars in the CFG-guided image is broken in the middle, which shows a misalignment with the text prompt. This type of artifact is substantially reduced in the CFG++ version, where the oars are depicted intact. The examples provided by SDXL in Fig. 4 further demonstrate the enhanced text-to-image alignment achieved with CFG++. For instance, the lion cub is not visible in Fig. 4(top-left), the dog appears with two tails in Fig. 4(top-right), the goggles have an unusual shape in Fig. 4(bottom-left), and the tree trunk is folded in Fig. 4(bottom-right). These artifacts, observed in the images generated by CFG, are absent in those produced by CFG++.
@@
In fact, this enhanced alignment capability is a natural consequence of CFG++, which directly minimizes the text-conditioned score-matching loss as shown in (12). In contrast, CFG indirectly seeks text alignment through the sharpened posterior distribution $p^{\omega}({\boldsymbol{x}}|{\boldsymbol{c}})\propto p({\boldsymbol{x}})p({\boldsymbol{c}}|{\boldsymbol{x}})^{\omega}$. Therefore, CFG++ inherently outperform CFG in terms of text alignment due to its fundamental design principle. For example, Fig. 5 displays the normalized text-conditioned score matching loss, represented as $(1-\bar{\alpha}_{t})\|{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})\|^{2}/\bar{\alpha}_{t}=\|{\boldsymbol{x}}-{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}\|^{2}$, throughout the reverse diffusion sampling process for both CFG and CFG++. The loss plot associated with CFG shows fluctuations and maintains a noticeable gap compared to CFG++ even after the completion of the reverse diffusion process. In fact, the fluctuation associated with CFG is also related to the off-manifold issue, and from Fig. 5 we can easily see that the off-manifold phenomenon is more dominant at early stage of reverse diffusion sampling. Conversely, the loss trajectory for CFG++ demonstrates a much smoother variation, particularly during the early stages of reverse diffusion.
@@
${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})$ $=({{\boldsymbol{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})})/{\sqrt{\bar{\alpha}_{t-1}}}$
$\simeq({{\boldsymbol{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})})/{\sqrt{\bar{\alpha}_{t-1}}}$ (22)
${\boldsymbol{x}}_{t}$ $=\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$
$\simeq\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})$ (23)

### v2

@@ 3.1 Derivation of CFG++
Instead of uncritically adopting the sharpened posterior distribution $p^{\omega}({\boldsymbol{x}}|{\boldsymbol{c}})\propto p({\boldsymbol{x}})p({\boldsymbol{c}}|{\boldsymbol{x}})^{\omega}$ as introduced by ho2021classifierfree, we adopt a fundamentally different strategy by reformulating text-guidance as an optimization problem. Specifically, our focus is on identifying a loss function $\ell({\boldsymbol{x}})$ in (8) such that, when minimized under the condition set by the text, enables the reverse diffusion process to generate samples that increasingly satisfy the text condition progressively.
One of the most significant contributions of this paper is to reveal that the text-conditioned score matching loss or alternatively, score distillation sampling (SDS) loss (poole2022dreamfusion) is ideally suited for our purpose. Specifically, we are interested in solving the following inverse problem through diffusion models:
@@
This implies that our goal is to identify solutions on the clean manifold ${\mathcal{M}}$ that optimally aligns with the text condition ${\boldsymbol{c}}$.
To avoid the Jacobian computation, in this paper, we attempt to solve (9) through DDS in (8). The resulting sampling process from reverse diffusion is then given by
${\boldsymbol{x}}_{t-1}$ $=$ $\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell_{sds}({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}.$ (10)
By using ${\boldsymbol{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\boldsymbol{x}}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}$ from the clean image ${\boldsymbol{x}}\in{\mathcal{M}}$, we can equivalently write the loss as $\ell_{sds}({\boldsymbol{x}})=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}\|{\boldsymbol{x}}-{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}\|^{2}$, which leads to
@@
where $\lambda:=\frac{2{\bar{\alpha}_{t}}}{{1-\bar{\alpha}_{t}}}\gamma_{t}$. Using the CFG notation ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}):={\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+{\lambda}[{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})]$ and ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}+\lambda({\hat{\boldsymbol{x}}_{\boldsymbol{c}}}-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})=({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$, (11) can be equivalently represented as
@@
which is summarized in Algorithm 2. By examining Algorithm 1 and Algorithm 2, we observe that CFG and CFG++ are mostly the same, with a crucial difference in the renoising process. This surprisingly simple fix of utilizing the unconditional noise ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ instead of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ leads to a smoother trajectory of generation, (Fig. 1 bottom) and generation with superior quality (Fig. 1 top).
Although we focus our construction of the solver on DDIM for simplicity, note that DDIM is just one way of reverse sampling. There are other widely-used solvers such as Karras Euler and its variants (karras2022elucidating), DPM-solver (lu2022dpm; lu2022dpmpp), their ancestral variants [^1] https://github.com/crowsonkb/k-diffusion, etc. For most widely used solvers up to the second order, a single-step update of solving the unconditional PF-ODE can be represented as
${\boldsymbol{x}}_{i}={{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{i-1})}+a_{i}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{i-1})+b_{i}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{i-2})+c_{i}{\boldsymbol{x}}_{i-1}+d_{i}{\boldsymbol{\epsilon}},\,{\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\bm{I}}),$ (14)
with $d_{i}\neq 0$ when one uses an ancestral sampler. Note that the first right-hand side term in (14) corresponds to the denosing, whereas the rest terms describes the higher-order corrected version of the renoising process. As the goal of CFG++ is to optimize the denoising process under the text-guidance while keeping the renoising components equivalent to unconditional sampling, applying CFG++ to the general iteration in (14) simply leads to
${\boldsymbol{x}}_{i}={\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{i-1})+a_{i}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{i-1})+b_{i}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{i-2})+c_{i}{\boldsymbol{x}}_{i-1}+d_{i}{\boldsymbol{\epsilon}},\,{\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\bm{I}}).$ (15)
Moreover, CFG++ naturally extends to distilled diffusion models, such as SDXL-turbo (sauer2023adversarial) and SDXL-lightning (lin2024sdxl), where akin to (15), we use the conditional denoised estimate, but for the rest of the noise components in the Euler solver, we use the unconditional estimate. For details in how to apply CFG++ to various solvers, see Appendix A.
@@ 3.2 Geometry of CFG++
Mitigating off-manifold phenomenon. In Fig. 1 (bottom), we illustrate the evolution of the posterior mean through Tweedie’s formula during the reverse diffusion process. Notably, in the early phases of reverse diffusion sampling under CFG, there is a sudden shift in the image and intense color saturation. Conversely, CFG++ is free from the undesirable off-manifold phenomenon. In the following, we investigate why this is the case.
Note that the denoised estimate of CFG and CFG++ at time $t$ can be equivalently represented as
@@
${\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})=(1-\lambda){\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\lambda{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t}),\quad{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})=(1-\omega){\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\omega{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t}).$ (16)
@@
$\lambda,\omega\in[0,1]$ facilitates an interpolation. However, with $\omega>1.0$, CFG extrapolates beyond the unconditional and conditional estimates. Consequently, under the assumption that the clean manifold be piecewise linear (chung2024decomposed), the conditional posterior mean estimates from CFG obtained with a guidance scale outside the range of $[0,1]$, can easily extend beyond the piecewise linear manifold. This may lead to the estimates potentially “falling off” the data manifold, as depicted by an orange arrow pointing downwards in Fig. 3(a). Thus, we select $\lambda\in[0,1]$ as the guidance scale for CFG++ to ensure it remains an interpolation between the unconditional and conditional estimate, thus preventing it from ‘falling off’ the clean data manifold. An additional source of the off-manifold phenomenon in CFG occurs during the transition from the clean manifold ${\mathcal{M}}$ to the subsequent noisy manifold ${\mathcal{M}}_{t-1}$, due to a similar off-manifold phenomenon from the large guidance scale (i.e. extrapolation), illustrated in Fig. 3 (b).
@@
Figure 4: Text-conditioned score matching loss throughout the reverse diffusion sampling for both CFG and CFG++ in SDXL. Avg. loss computed with 55 prompts from (chen2024pixartalpha).
@@
Text-Image alignment. In Fig. 1 (top), and Fig. 9, we observe enhanced text-to-image alignment achieved with CFG++. This enhanced alignment capability is a natural consequence of CFG++, which directly minimizes the text-conditioned score-matching loss as shown in (9). In contrast, CFG indirectly seeks text alignment through the sharpened posterior distribution $p^{\omega}({\boldsymbol{x}}|{\boldsymbol{c}})\propto p({\boldsymbol{x}})p({\boldsymbol{c}}|{\boldsymbol{x}})^{\omega}$. Therefore, CFG++ inherently outperforms CFG in terms of text alignment due to its fundamental design principle. For example, Fig. 4 displays the normalized text-conditioned score matching loss, represented as $(1-\bar{\alpha}_{t})\|{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})\|^{2}/\bar{\alpha}_{t}=\|{\boldsymbol{x}}-{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}\|^{2}$, throughout the reverse diffusion sampling process for both CFG and CFG++. The loss plot associated with CFG shows fluctuations and maintains a noticeable gap compared to CFG++ even after the completion of the reverse diffusion process. In fact, the fluctuation associated with CFG is also related to the off-manifold issue, and from Fig. 4 we can easily see that the off-manifold phenomenon is more dominant at early stage of reverse diffusion sampling. Conversely, the loss trajectory for CFG++ demonstrates a much smoother variation, particularly during the early stages of reverse diffusion.
@@
${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})$ $=({{\boldsymbol{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})})/{\sqrt{\bar{\alpha}_{t-1}}}\simeq({{\boldsymbol{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})})/{\sqrt{\bar{\alpha}_{t-1}}}$ (17)
${\boldsymbol{x}}_{t}$ $=\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})\simeq\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})$ (18)